# Quantum accessible information and classical entropy inequalities

### v1

@@ Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states [2] is reconsidered and the corresponding entropy inequalities are proposed. We prove these inequalities in the cases of state ensembles corresponding to acute or flat pyramids, thus providing the first proof of the hypotheses concerning globally information-optimal observables put forward in the paper [2].
@@ Introduction
The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, and important in particular for applications in quantum cryptography, see e.g. the survey [14]. The problem was solved in some special (notably, symmetric) cases but in general remains open. In this paper we describe a general approach following [7] which gives a criterion for optimality and ties up the problem with entropy inequalities – the tight lower bounds for the classical Shannon information. It turns out that the optimality criterion [7] when applied to specific ensembles of states leads to new nontrivial entropy inequalities that are discrete “distant relatives” of the famous log-Sobolev inequality. In this light, the hypothesis of an information-optimal measurement for an ensemble of equiangular equiprobable states – quantum pyramids formulated and numerically substantiated by Englert and Řeháček in [2] – is reconsidered in the present paper. In the case of acute pyramids the corresponding parametric family of entropy inequalities (23) is proposed in the section 4.1. The inequalities suggest the new tight lower bounds for the Shannon entropy of a discrete probability distribution as compared e.g. to the known bound from [11], theorem 3.11.
The proof of the inequality (23), given in section 5, was obtained by the second author as a response to the conjecture put forward by the first author. Via the optimality criterion of section 2, it also implies the validity of a hypothesis concerning accessible information for state ensembles of acute quantum pyramids, formulated and numerically substantiated in [2] (corollary 1 to theorem 2).
In the section 4.4 the case of obtuse pyramids is considered to get further insight and to formulate the hypotheses concerning the corresponding entropy inequalities (60), (64). Our approach allows also to derive the nonlinear equations (50) (resp. (62)) serving to determine the critical values of the parameter $t$ of the optimal observable corresponding to transition from moderately to strongly acute (resp, obtuse) pyramids.
In the section 4.5 we derive and prove the tight entropy inequality (66) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 2). The findings in sections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and is the optimal one in the flat case (tentatively also in the obtuse case).
@@
It was shown [1] that the supremum is attained on an observable of the form $M_{k}=\left|\varphi_{k}\rangle\langle\varphi_{k}\right|,\,k=1,\dots,n$ with $d\leq n\leq d^{2}$ and linearly independent $M_{k}$. In particular, the supremum can be replaced with the maximum.
@@
$M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\pi_{j}\rho_{j}\right)\overline{\rho}^{-1/2}$
@@
$A(\mathcal{E})=H(\pi)-\mathrm{Tr\,}\overline{\rho}\Lambda_{0}.$
@@ 4.1 The lower bound for the Shannon entropy: formulation and discussion
The inequality (23) arises in connection with the problem of accessible information for ensemble of $m$ equiprobable equiangular pure states discussed in the next subsection. The proof of (23) will be given in section 5. Note that in the case $m=2$ the inequality (23) reduces to (19).
@@ 4.2 The conjecture: “quantum pyramids”
$\bar{\rho}\equiv\frac{1}{m}\sum_{k\neq j}^{m}|\psi_{j}\rangle\langle\psi_{j}|=r_{1}I+\left(r_{0}-r_{1}\right)|e_{0}\rangle\langle e_{0}|.$ (34)
@@
In what follows we consider the solution for the accessible information $A(\mathcal{E})$ and its optimizing observable conjectured and supported numerically in [2]. In the rest of this section and in the next section we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form =$\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
@@
$A(\mathcal{E})=\left\{\begin{array}[]{cc}I^{(SRM)},&r_{0}\leq\frac{4(m-1)}{m-2}\\ \frac{m\left(1-r_{0}\right)}{m-2}\log\sqrt{\frac{r_{1}}{r_{0}}},&r_{0}>\frac{4(m-1)}{m-2}\end{array}\right.,$ (39)
@@
-- $\frac{m-1}{m}\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}$
@@ 4.3 Checking the optimality conditions: the case of acute pyramids
Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 in previous section we look for $\Lambda_{0}$ in the form
@@
the values which do not depend on the parameters of the pyramid $r_{0},r_{1}.$ The criterion for the applicability of this second case is $t_{a}(m)\equiv\sqrt{\frac{r_{1}}{r_{0}}}\tau_{a}(m)=\sqrt{\frac{r_{1}}{r_{0}}}\frac{2(m-1)}{m-2}<1$ which, via (30), is equivalent to $r_{0}>\frac{4(m-1)}{m^{2}}$, and, via (33), to $a^{2}<\frac{m-1}{m}.$
Correspondingly, the condition for applicability of the solution (48) is $a^{2}\geq\frac{m-1}{m}.$ By introducing the parameter $p=a^{2}$ we have $b=\sqrt{\frac{1-p}{m-1}},$ leading to
@@
Finally, from (46) and (34)
$\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=c_{1}+c_{0}r_{0}$
and
$A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\log m+p\log p+\left(1-p\right)\log\frac{1-p}{m-1}$ (53)
agrees with the expression (39). Thus we have the following corollary to the theorem 2:
@@
For ensemble of of acute pyramids the information-optimal observable is SRM given by the equations (37), (38) as conjectured in [2]. The accessible information is given by the expression (53).
@@ 4.4 The case of obtuse pyramids
The generalization of this picture to the case $m\geq 3$ based on extensive numerical studies was suggested and elaborated in [2]. Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight. We consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (29) with $r_{0}<r_{1}$ $(b<0)$ The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
@@
$|\tilde{\phi}_{k}\rangle$ $=$ $t^{-1}\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=t^{-1}\left(\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle\right);$
$p_{k}$ $=$ $t^{-2}\left(r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}\right);\quad k=1,\dots,m;$
@@
$\tilde{\mu}_{0}(p)=\frac{\sqrt{\frac{p\left(1-p\right)}{m-1}}\left(\log p-\log\frac{1-p}{m-1}\right)}{\left(\left(m-1\right)\sqrt{\frac{1-p}{m-1}}-\sqrt{p}\right)\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)},$
$\tilde{\mu}_{1}(p)=\frac{\sqrt{p}\log p+\sqrt{\frac{1-p}{m-1}}\log\frac{1-p}{m-1}}{\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)}.$
@@
The numerically found values:
@@
Here $mr_{0}(m)$ show a good agreement with the values found empirically for $m\leq 6$ in [2], eq. (39). In the case $m=7$, as well as for greater values of $m$, the solution of the equation (62) becomes negative and also $p(m)<\frac{m-1}{m}.$ This makes plausible the hypothesis that for $m=7,\dots$ the optimal observable is SRM, and the corresponding entropy inequality is (60) for all $p\geq\frac{m-1}{m}.$
@@
The equality is attained if $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$ for $j\geq 3$ and for all permutations of these $z_{j}.$
In this case the accessible information is
$A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\log\frac{m}{2}-r_{0}\left(\tau_{o}(m)^{-1}-1\right)\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}.$
@@ 4.5 The case of flat pyramid
The theorem 1 then implies
@@
The corresponding accessible information is [^5] Cf. [2], second line in eq. (33).
@@
5 Proof of theorem 2
Let $m\geq 2,\ p\geq\dfrac{m-1}{m}$ and let $\mu_{0}(p),\mu_{1}(p)$ be given by (24), (25).
For $p\in[(m-1)/m,1]$ and a probability distribution $(t_{1},...,t_{m})$, it holds
@@ 5 Proof of theorem 2
$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}.$ (71)
@@
Let $A\geq 0,\ B,\ C\in\mathbb{R}$. Then the function
@@
reaches a local minimum on $[0,1]$ at no more than two points $t_{1}\leq 1/2\leq t_{2}$ which are critical. In this case, if $B>0$ and $t_{*}>1/2$ is critical ($B<0$ and $t_{*}<1/2$), then $t_{*}$ is a point of the global minimum.
@@
$\alpha\in[0,1/2],\quad p\in[1-\alpha,1],\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
@@
      $\ln\dfrac{1-\alpha}{\alpha}\geq 2-4\alpha,\quad\alpha\in(0,1/2].$
@@
    It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$. Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$, and $f(t)\geq 0$.

### v2

@@ Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states [2] is reconsidered and the corresponding entropy inequalities are proposed and proved. Via the optimality criterion, this provides also the first proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
@@ Introduction
The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, and important in particular for applications in quantum cryptography, see e.g. the survey [14]. The problem was solved in some special (notably, symmetric) cases but in general remains open. In section (2) of this paper we describe a general approach following [7] which gives a criterion for optimality and ties up the problem with entropy inequalities – the tight lower bounds for the classical Shannon information. It turns out that the optimality criterion when applied to specific ensembles of states leads to new nontrivial entropy inequalities that are discrete “distant relatives” of the famous log-Sobolev inequality, as explained in section 3.
In this light, the conjecture of an information-optimal measurement for an ensemble of equiprobable equiangular states – quantum pyramid, formulated and numerically substantiated by Englert and Řeháček in [2] – is reconsidered in section 4 of the present paper. In the case of acute pyramids the corresponding parametric family of entropy inequalities (27) is proposed in subsections 4.1-4.3. The inequalities suggest the new tight lower bounds for the Shannon entropy of a discrete probability distribution as compared e.g. to the known bound from [11], theorem 3.11.
In the subsection 4.4 the case of obtuse pyramids is considered to get further insight and to formulate the hypotheses concerning the corresponding entropy inequalities (66), (72). Our approach allows also to derive the nonlinear equations (54) (resp. (70)) serving to determine the critical values of the parameter $t$ of the optimal observable corresponding to transition from moderately to strongly acute (resp, obtuse) pyramids.
In the subsection 4.5 we derive and prove the tight entropy inequality (75) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 3) in the case of flat pyramids. The findings in subsections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and is the optimal one.
The proofs of the inequality (27), given in subsection 5.1, as well as of the inequalities (66), (72) in subsection 5.2 was obtained by the second author as a response to the hypotheses put forward by the first author. Via the optimality criterion of section 2, it also implies the validity of the conjectures concerning accessible information for quantum pyramids, formulated and numerically substantiated in [2] (corollary 1 to theorem 2, corollary 2 to theorem 3).
@@
It was shown [1] that the supremum is attained on an observable of the form $M_{k}=\left|\varphi_{k}\rangle\langle\varphi_{k}\right|,\,k=1,\dots,n$ with $d\leq n\leq d^{2}$ and linearly independent $M_{k}$ (if the matrices of $\rho_{j}$ in some basis are real then the number of components is reduced to $\frac{d(d+1)}{2}$). In particular, the supremum can be replaced with the maximum.
@@
$M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\pi_{j}\rho_{j}\right)\overline{\rho}^{-1/2}$ (1)
@@
$A(\mathcal{E})=H(\pi)-\mathrm{Tr\,}\overline{\rho}\Lambda_{0}.$ (9)
@@
Especially important for us is the case of pure-states ensemble with $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ where $|\psi_{j}\rangle;j=1,\dots,d$, are linearly independent vectors. In this case the resolution of the identity (1) is
$M_{j}^{\prime}=|e_{j}\rangle\langle e_{j}|,$
where the vectors
$|e_{j}\rangle=\sqrt{\pi_{j}}\bar{\rho}^{-1/2}|\psi_{j}\rangle;\quad j=1,\dots,d,$ (15)
by necessity form an orthonormal basis. Then the entropy inequality (7) takes the form
$-\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (16)
where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint $\sum_{j=1}^{d}\left|z_{j}\right|^{2}=1$. The left-hand side is just the Shannon entropy of the p.d. $\left\{\left|z_{j}\right|^{2};j=1,\dots,d\right\}$, and such an inequality can be approached with the tools of classical analysis.
@@ 4.1 The lower bound for the Shannon entropy: formulation and discussion
The inequality (27) arises in connection with the problem of accessible information for ensemble of $m$ equiprobable equiangular pure states discussed in the next subsection. The proof of (27) will be given in section 5.1. Note that in the case $m=2$ the inequality (27) reduces to (23).
@@ 4.2 The conjecture: “quantum pyramids”
$\bar{\rho}\equiv\frac{1}{m}\sum_{j=1}^{m}|\psi_{j}\rangle\langle\psi_{j}|=r_{1}I+\left(r_{0}-r_{1}\right)|e_{0}\rangle\langle e_{0}|.$ (38)
@@
In what follows we consider the solution for the accessible information $A(\mathcal{E})$ and its optimizing observable conjectured and supported numerically in [2]. In the rest of this subsection and in the next subsection we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form =$\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
@@
$A(\mathcal{E})=\left\{\begin{array}[]{cc}I^{(SRM)},&r_{0}\leq\frac{4(m-1)}{m^{2}}\\ \frac{m\left(1-r_{0}\right)}{m-2}\log(m-1),&r_{0}>\frac{4(m-1)}{m^{2}}\end{array}\right.,$ (43)
@@
$+$ $\frac{m-1}{m}\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}$
@@ 4.3 Checking the optimality conditions: the case of acute pyramids
Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 in previous subsection we look for $\Lambda_{0}$ in the form
@@
the values which do not depend on the parameters of the pyramid $r_{0},r_{1}.$ The criterion for the applicability of this second case is
$t_{a}(m)\equiv\sqrt{\frac{r_{1}}{r_{0}}}\tau_{a}(m)=\sqrt{\frac{r_{1}}{r_{0}}}\frac{2(m-1)}{m-2}<1$
which, via (34), is equivalent to $r_{0}>\frac{4(m-1)}{m^{2}}$, and, via (37), to $a^{2}<\frac{m-1}{m}.$
Correspondingly, the condition for applicability of the solution (52) is $a^{2}\geq\frac{m-1}{m}.$ By introducing the parameter $p$ such that
$p=a^{2},\quad\frac{1-p}{m-1}=b^{2},$ (56)
we have $b=\sqrt{\frac{1-p}{m-1}},$ leading to
@@
Finally, from (50) and (38) $\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=c_{1}+c_{0}r_{0}$ so that
$A(\mathcal{E})=\log m-c_{1}-c_{0}r_{0}.$ (58)
Thus we have the following corollary to the theorem 2:
@@
For ensemble of acute pyramids the information-optimal observable is given by the equations (41), (42) as conjectured in [2]. The accessible information is given by the expressions
$A(\mathcal{E})=\left\{\begin{array}[]{cc}\log m+p\log p+\left(1-p\right)\log\frac{1-p}{m-1},&p\in\left[\frac{m-1}{m},1\right]\\ \frac{m-1}{m-2}\log(m-1)\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)^{2},&p\in\left[\frac{1}{m},\frac{m-1}{m}\right]\end{array}\right.$ (59)
which agree with (43).
Proof.
Taking into account theorem 2, the first statement was in fact proved above in this section. For accessible information we use the relation (58). A routine calculation using (52), (37), (56) gives (45) which is the first line in (59) for moderately acute pyramids. For strongly acute pyramid, using (55), one gets
$A(\mathcal{E})=\frac{m\left(1-r_{0}\right)}{m-2}\log(m-1),$
which is the second line, taking into account that
$1-r_{0})=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}$
by (34), (37). ∎
@@ 4.4 The case of obtuse pyramids
The generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2]. Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight. We consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (33) with $r_{0}<r_{1}$ $(b<0)$ The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
@@
$|\tilde{\phi}_{k}\rangle$ $=$ $t^{-1}\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=t^{-1}\left(\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle\right),$
$p_{k}$ $=$ $t^{-2}\left(r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}\right);\quad k=1,\dots,m,$
@@
$\tilde{\mu}_{0}(p)=\frac{\sqrt{\frac{p\left(1-p\right)}{m-1}}\left(\log p-\log\frac{1-p}{m-1}\right)}{\left(\left(m-1\right)\sqrt{\frac{1-p}{m-1}}-\sqrt{p}\right)\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)},$ (67)
$\tilde{\mu}_{1}(p)=\frac{\sqrt{p}\log p+\sqrt{\frac{1-p}{m-1}}\log\frac{1-p}{m-1}}{\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)}.$ (68)
@@
The numerically found values [^5] By using equation (70) it is possible to express the function $p(m)$ via the Lambert $W$-function.:
@@
Here $mr_{0}(m)$ show a good agreement with the values found empirically for $m\leq 6$ in [2], eq. (39). In the case $m=7$, as well as for greater values of $m$, the solution of the equation (70) becomes negative and also $p(m)<\frac{m-1}{m}.$ This makes plausible the hypothesis that for $m\geq 7$ the optimal observable is SRM, and the corresponding entropy inequality is (66) for all $p\geq\frac{m-1}{m}.$
@@
The equality is attained for $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$, $j\geq 3$; for $z_{1}=\sqrt{p(m)},\,z_{j}=-\sqrt{\frac{1-p(m)}{m-1}},\,j\geq 2$; and for all permutations of these $z_{j}.$
In this case the accessible information, by using (58), is
$A(\mathcal{E})$ $=$ $\log\frac{m}{2}-r_{0}\left(\tau_{o}(m)^{-1}-1\right)\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}$
$=$ $\log\dfrac{m}{2}-r_{0}m\tilde{\mu}_{0}(p(m)).$
The following theorem summarizes the content of this section.
Theorem 3.
If $m\geq 7$ then $p(m)<\frac{m-1}{m}$ and the entropy inequality (66) holds for $p\in\left[\frac{m-1}{m},1\right]$. If $m\leq 6$ then $p(m)>\frac{m-1}{m}$ and the inequality (66) holds for $p\in\left[p(m),1\right]$, yielding the inequality (72) for $p\in\left[\frac{m-1}{m},p(m)\right]$.
The proof of the inequalities is given in section 5.2.
Corollary 2.
In the case of obtuse pyramids the information-optimal observable is given by the relations (61) with $t=1$ for $m\geq 7$, and $t=\sqrt{\frac{r_{1}}{r_{0}}}\tau_{o}(m)$ for $m\leq 6$. The accessible information is given by the expression (45) if $p\in[\max\left\{\frac{m-1}{m},p(m)\right\},1],$ otherwise by (4.4) if $m\leq 6$ and $p\in\left[\frac{m-1}{m},p(m)\right]$.
@@ 4.5 The case of flat pyramid
Theorem 1 then implies
@@
The corresponding accessible information is
@@
5 Proofs
5.1 Proof of theorem 2
Let $m\geq 2,\ \dfrac{m-1}{m}\leq p<1$ and let $\mu_{0}(p),\mu_{1}(p)$ be given by (28), (29).
For $p\in[(m-1)/m,1)$ and a probability distribution $(t_{1},...,t_{m})$, it holds
@@ 5.1 Proof of theorem 2
In the special case $p=1$, the inequality (78) is trivially satisfied by substituting $\mu_{0}(1):=0,\ \mu_{1}(1)=0$.
@@
$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$ (80)
for real variables $p$ and $\alpha$ varying in the range $(0,1)$.
@@
Let $A>0,\ B,\ C\in\mathbb{R}$. Then the function
@@
[…]
@@
$\alpha\in{(0,1/2],\quad p\in[1-\alpha,1},\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
@@
      $\ln\dfrac{1-\alpha}{\alpha}\geq 2-4\alpha.$
@@
    It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$. Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$. We obtained the desired inequality $f(t)\geq 0$ for the minimizer $t_{*}$ and it is even more so for the other points $t\in[0,1]$.